# Investigating the Robustness of Extreme Precipitation Super-Resolution Across Climates

### v2

@@ Abstract
The coarse spatial resolution of gridded climate models, such as general circulation models, limits their direct use in projecting socially relevant variables like extreme precipitation. Most downscaling methods estimate the conditional distributions of extremes by generating large ensembles, complicating the assessment of robustness under distributional shifts, such as those induced by climate change. To better understand and potentially improve robustness, we propose super-resolving the parameters of the target variable’s probability distribution directly using analytically tractable mappings. Within a perfect-model framework over Switzerland, we demonstrate that vector generalized linear and additive models can super-resolve the generalized extreme value distribution of summer hourly precipitation extremes from coarse precipitation fields and topography. We introduce the notion of a “robustness gap”, defined as the difference in predictive error between present-trained and future-trained models, and use it to diagnose how model structure affects the generalization of each quantile to a pseudo-global warming scenario. By evaluating multiple model configurations, we also identify an upper limit on the super-resolution factor based on the spatial auto- and cross-correlation of precipitation and elevation, beyond which coarse precipitation loses predictive value. Our framework is broadly applicable to variables governed by parametric distributions and offers a model-agnostic diagnostic for understanding when and why empirical downscaling generalizes to climate change and extremes.
@@
[1]organization=Expertise Center for Climate Extremes, University of Lausanne, addressline=, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@
[2]organization=Federal Office of Meteorology and Climatology MeteoSwiss, addressline=, city=Zurich, postcode=8058, state=ZH, country=Switzerland
@@
[3]organization=Faculty of Geosciences and Environment, University of Lausanne, addressline=, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@
[4]organization=Faculty of Business and Economics (HEC), University of Lausanne, addressline=, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@ 1 Introduction
Climate adaptation requires projecting high-impact weather events at local scales, notably extreme precipitation due to its impacts on ecosystems and infrastructure (Fowler et al., 2007; Gimeno et al., 2022). General circulation models (GCMs) are too coarse (with horizontal grid spacing around 100 km) to explicitly simulate such extremes (Maraun, 2016; Benestad, 2004). This limitation warrants the use of expensive regional climate models that can only be run selectively, especially at convection-permitting scales (Schär et al., 2020), potentially under-sampling projection uncertainty (Hawkins and Sutton, 2011). Empirical downscaling, including statistical (Maraun and Widmann, 2018) and machine learning (ML; Rampal et al. (2024b)) methods, offers a promising complement or alternative by directly predicting relevant variables at local scales when suitable training data are available. Recent advances in ML-based super-resolution (Wang et al., 2022) and generative modeling (Yang et al., 2023) have further fueled the rapid development of empirical downscaling for precipitation (Rampal et al., 2025; Srivastava et al., 2024; Rampal et al., 2022; Vandal et al., 2019). However, their application to climate change remains limited by challenges in understanding their robustness—i.e., how well they extrapolate to warmer climates (Hernanz et al., 2022). This is especially true for extremes, where stationarity is difficult to assess from historical performance alone (Dixon et al., 2016).
@@
This motivates pseudo-reality experiments (also called “model as truth” or “perfect model”), in which outputs from a dynamical regional climate model are treated as pseudo-observations for empirical downscaling, enabling direct benchmarking of generalization capabilities (Maraun et al., 2015). While such experiments have helped identify best practices (e.g., optimal predictor sets) for improving the robustness of simple statistical downscaling algorithms (Charles et al., 1999; Vrac et al., 2007; Dayon et al., 2015), they remain under-used for more sophisticated methods. Notable exceptions include Legasa et al. (2023), who showed that a posteriori random forests targeting parameters of the precipitation gamma distribution (Legasa et al., 2022) generalize better to a warmer climate than generalized linear models and convolutional neural networks; Baño-Medina et al. (2024), who found that deep learning emulators generalize more reliably when GCM predictors are bias-adjusted to the upscaled regional model; and Rampal et al. (2024a), who showed that generative adversarial networks outperform deterministic convolutional neural networks in projecting warming-driven precipitation extremes. An emerging challenge is that these results remain scattered—typically evaluated per case or quantile—and we still lack a simple framework to explain what drives robustness across varying degrees of extremeness.
@@
Here, we aim to address this gap for the task of super-resolving precipitation extremes, i.e., improving their spatial resolution. We focus on this task because it is tractable: (1) although fine-resolution and coarse-resolution extremes do not occur simultaneously, they typically fall within the same temporal block, making them weakly paired; and (2) super-resolution is generally not treated on a per-GCM basis. This contrasts with the full empirical downscaling pipeline required for local climate projections, where global and sometimes even regional climate model outputs must be bias-corrected (François et al., 2020; Cannon, 2018; Vrac and Friederichs, 2015) before being brought to the desired spatial scale. We emphasize that super-resolution is not trivial, especially for precipitation extremes, where analog-based and deterministic methods often fail (Sachindra et al., 2018; Reddy et al., 2023). We therefore introduce the concept of super-resolving extremal distributions, where we learn to increase the spatial resolution of a distribution’s parameters. Friederichs (2010) previously downscaled precipitation block maxima by targeting GEV parameters from large-scale circulation features. However, we focus here on cases where convection-permitting extremes (1–5 km) are at least partially resolved at coarser resolutions (12–50 km), consistent with the HighResMIP2 range targeted by next-generation of Earth system models (Roberts et al., 2025), allowing a super-resolution setup.
@@
As illustrated in Figure 1, our generative super-resolution model combines regression-based modeling with extreme value theory (Vrac and Naveau, 2007). It learns the parameters of the Generalized Extreme-Value (GEV) distribution governing fin-resolution precipitation maxima from those governing their coarse-resolution counterparts. To ensure interpretability, we use Vector Generalized Additive Models (VGAM) and Vector Generalized Linear Models (VGLM), which allow us to specify a distribution family and examine how features influence predicted distributions through spline terms. This approach offers three key advantages: low training costs, reliable probability estimates, and transparent model behavior.
@@
To quantify our models’ robustness to climate change, we adopt a pseudo-reality framework over Switzerland, where complex topography—ranging from alpine areas above 4,000 m to low-lying regions below 200 m—drives strong spatial variability in precipitation. Observations show a positive trend in annual maximum daily precipitation at 91% of Swiss stations, with 10-minute summer precipitation intensities increasing at 5.7% per decade (Bauer and Scherrer, 2024; Scherrer et al., 2016). Convection-permitting model ensembles reproduce these changes more faithfully than coarser models and project a 6-7% increase in heavy summer precipitation intensity per degree of warming, despite an overall summer mean decrease (Estermann et al., 2025; Ban et al., 2021). As summer precipitation is dominated by convective events, it remains particularly challenging to downscale without underestimating variability (Zubler et al., 2014; Schmidli et al., 2007), though this can be partly mitigated by explicitly incorporating temperature dependence in precipitation scaling (Moraga et al., 2024).
@@
Motivated by the challenge of super-resolving extreme summer precipitation, we base our pseudo-reality experiment on the present-future convection-permitting simulation pair of Hentgen et al. (2019) (Section 2). We then formalize the task of super-resolving distributions and define the “robustness gap” to quantitatively assess model generalization across climates (Section 3). Applied to the super-resolution of the GEV using VGAMs and VGLMs (Section 4), our diagnostic framework explains generalization errors across quantiles, interprets model behavior via splines, and identifies performance limits when the super-resolution factor becomes too large (Section 5). We conclude in Section 6. The Supplementary Material (SM) provides technical derivations that support Sections 4 and 5 of the manuscript.
@@ 2 Data
We super-resolve hourly precipitation model output over Switzerland for 11 European summer (JJA) seasons in both historical (1999–2009) and projected (2079–2089) climates. Swiss summers are characterized by weak synoptic forcing, frequent afternoon thunderstorms and convective rainfall. Data are derived from simulations using the regional weather and climate model COSMO (Consortium for Small-scale Modeling), run over a European domain of approximately $3000\times 3000$ km at 2.2 km grid spacing. Full simulation details are available in Hentgen et al. (2019), with a brief overview provided in Section 2.1. Section 2.2 describes the construction of the coarse-resolution field from the fine-resolution data. The elevation statistics employed are detailed in Section 2.3, and the training, validation, and test set configurations are outlined in Section 2.4.
@@ 2.1 Simulations
We calibrate and test our method using regional climate simulations at 2.2 km horizontal resolution. Kilometer-scale climate modeling has become increasingly common in recent years (Ban et al., 2021; Stevens et al., 2019), offering several advantages over coarser-resolution simulations. The finer representation of topography and land surface enables more realistic precipitation patterns in complex terrain such as the Alps. In addition, vertical air motion is explicitly resolved by the governing equations, bringing the model formulation closer to physical first principles. Compared to coarser resolutions that require convective parameterization, explicit convection improves the realism of the hydrological cycle, particularly for extreme precipitation and its associated mechanisms and feedbacks (Prein et al., 2015; Schär et al., 2020; Lenderink et al., 2025).
To assess generalization, we use model output from a simulation based on the pseudo-global warming (PGW) method (Rasmussen et al., 2011; Adachi and Tomita, 2020; Schär et al., 1996). PGW aims to simulate a warmer climate by preserving the spatiotemporal structure of historical weather patterns. A regional simulation is first performed using reanalysis-based boundary conditions, referred to as the control simulation (CTRL). This serves as the baseline for comparison with the PGW scenario. A physically consistent climate change signal (“climate delta”) is then applied to these boundaries, and the simulation is repeated. The deltas are derived from multi-year climatological means, representing differences between the MPI-ESM-LR global climate model’s future (RCP 8.5) and historical scenarios (Kröner et al., 2017), and vary with latitude, longitude, elevation, and month. The PGW simulation hence retains the sequence of weather events from CTRL, adjusted by the imposed climate signal. As such, the PGW simulation primarily captures thermodynamic changes (e.g., increases in temperature and moisture) while preserving large-scale circulation patterns from the historical record (Hall et al., 2024). This reduces confounding effects from circulation biases, providing a dataset of intermediate complexity with greater control for evaluating the generalization of statistical downscaling methods to warmer climates.
@@ 2.2 Coarse-Graining
To ensure comparability across climates, we adopt an idealized super-resolution framework in which the coarse-resolution inputs are coarsened versions of the fine-resolution targets. Hourly precipitation data are aggregated to coarser grid spacings via mean pooling, i.e., using a square spatial filter of prescribed length on the native COSMO grid. To assess robustness across not only climates but also super-resolution factors, we apply pooling lengths of 13.2 km, 26.4 km, and 52.8 km. We then compute monthly maxima on the coarsened low-resolution data (features) and on the fine-resolution data (targets).
@@
As shown in Figure 2, decreasing spatial resolution leads to substantial information loss. Precipitation values become more uniform, with reduced means and variances of extremes consistent with the geostatistical concept of “change of support” (Park, 2013; Onibon et al., 2004). The highest extremes are particularly affected, with distribution tails shortening as resolution decreases. This is due to the localized nature of extremes: coarse-resolution cells blend multiple events within a grid block, smoothing out their intensity and variability. These patterns remain consistent across all resolutions in both climates. Future climates exhibit broader, flatter distributions with heavier tails and higher standard deviations, indicating increased dispersion.
@@ 2.4 Training, Validation, and Test Split
To prevent overfitting and to ensure objective model evaluation, we define spatially separated splits (Valavi et al., 2019; Brenning, 2012), shown in Figure 3. Rather than assigning entire contiguous regions to each split, we partition the domain into 10 spatial regions, which are then distributed across the training, validation, and test sets. This approach preserves spatial separation while maintaining representativeness of the overall data distribution in each set. The training set (white) includes 70% of the data and is used to optimize the model’s trainable parameters, providing enough information to learn the super-resolution mapping. The validation set (dark gray), comprising 17% of the data, is used for model and feature selection and helps prevent overfitting. The remaining 13% forms the test set (light gray), which evaluates the model’s ability to generalize to unseen regions. We refer to the generalization quantified by the test set as “spatial generalizability” to distinguish it from the models’ ability to generalize across climates.
@@ 3.1 Super-Resolution of Distributions
A concrete example for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of regular two-dimensional grids with mesh sizes $\delta_{1}>0$ and $\delta_{2}>0$, each comprising $m$ points along each axis:
$\mathcal{G}_{m}^{\delta_{1}}\overset{\mathrm{def}}{=}\left\{(\delta_{1}\cdot i_{1},\delta_{1}\cdot i_{2}):(i_{1},i_{2})\in\llbracket 1,m\rrbracket^{2}\right\},$ (2)
$\mathcal{G}_{m}^{\delta_{2}}\overset{\mathrm{def}}{=}\left\{(\delta_{2}\cdot i_{1},\delta_{2}\cdot i_{2}):(i_{1},i_{2})\in\llbracket 1,m\rrbracket^{2}\right\},$ (3)
where $\delta_{1}>\delta_{2}$, so that $\mathcal{G}_{m}^{\delta_{2}}$ has higher spatial resolution than $\mathcal{G}_{m}^{\delta_{1}}$.
@@
where $\bm{\tau}_{\bm{s}}$ represents a set of features specific to grid point $\bm{s}$ (e.g., topographic or land-use characteristics), and $f$ is a function—possibly learned from data—that maps coarse-resolution distributional information and local covariates to fine-resolution distribution parameters.

### v3

@@ Abstract
The coarse spatial resolution of gridded climate models, such as general circulation models, limits their direct use in projecting socially relevant variables like extreme precipitation. Most downscaling methods estimate the conditional distributions of extremes by generating large ensembles, complicating the assessment of robustness under distributional transformations, such as those induced by climate change. To better understand and potentially improve robustness, we propose super-resolving the parameters of the target variable’s probability distribution directly using analytically tractable mappings. Within a perfect-model framework over Switzerland, we demonstrate that vector generalized linear and additive models can super-resolve the generalized extreme value distribution of summer hourly precipitation extremes from coarse precipitation fields and topography. We introduce the notion of a “robustness gap”, defined as the difference in predictive error between present-trained and future-trained models, and use it to diagnose how model structure affects the generalization of each quantile to a pseudo-global warming scenario. By evaluating multiple model configurations, we also identify an upper limit on the super-resolution factor based on the spatial auto- and cross-correlation of precipitation and elevation, beyond which coarse precipitation loses predictive value. Our framework is broadly applicable to variables governed by parametric distributions and offers a model-agnostic diagnostic for understanding when and why empirical downscaling generalizes to climate change and extremes.
@@
[1]organization=Expertise Center for Climate Extremes, University of Lausanne, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@
[2]organization=Federal Office of Meteorology and Climatology MeteoSwiss, city=Zurich, postcode=8058, state=ZH, country=Switzerland
@@
[3]organization=Faculty of Geosciences and Environment, University of Lausanne, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@
[4]organization=Faculty of Business and Economics (HEC), University of Lausanne, city=Lausanne, postcode=1015, state=VD, country=Switzerland
@@ 1 Introduction
Climate adaptation requires projecting high-impact weather events at local scales, notably extreme precipitation due to its impacts on ecosystems and infrastructure (fowler2007; gimeno22). General circulation models (GCMs) are too coarse (with horizontal grid spacing around 100 km) to explicitly simulate such extremes (maraun2016bias; benestad04). This limitation warrants the use of expensive regional climate models that can only be run selectively, especially at convection-permitting scales (schar20), potentially under-sampling projection uncertainty (hawkins2011potential). Empirical downscaling, including statistical (maraun2018statistical) and machine learning (ML; rampal2024enhancing) methods, offers a promising complement or alternative by directly predicting relevant variables at local scales when suitable training data are available. Recent advances in ML-based super-resolution (wang2022comprehensive) and generative modeling (yang2023diffusion) have further fueled the rapid development of empirical downscaling for precipitation (rampal2024reliableGAN; srivastava2024precipitation; Rampal2022interpretabledeeplearning; vandal2019intercomparison). However, their application to climate change remains limited by challenges in understanding their robustness—i.e., how well they extrapolate to warmer climates (hernanz22suitability). This is especially true for extremes, where stationarity is difficult to assess from historical performance alone (dixon2016evaluating).
@@
This motivates pseudo-reality experiments (also called “model as truth” or “perfect model”), in which outputs from a dynamical regional climate model are treated as pseudo-observations for empirical downscaling, enabling direct benchmarking of generalization capabilities (maraun2015value). While such experiments have helped identify best practices (e.g., optimal predictor sets) for improving the robustness of simple statistical downscaling algorithms (charles1999validation; vrac2007general; dayon2015transferability), they remain under-used for more sophisticated methods. Notable exceptions include legasa23, who showed that a posteriori random forests targeting parameters of the precipitation gamma distribution (legasa22aposterioriRF) generalize better to a warmer climate than generalized linear models and convolutional neural networks; bano24transferability, who found that deep learning emulators generalize more reliably when GCM predictors are bias-adjusted to the upscaled regional model; and rampal2024extrapolation, who showed that generative adversarial networks outperform deterministic convolutional neural networks in projecting warming-driven precipitation extremes. An emerging challenge is that these results remain scattered—typically evaluated per case or quantile—and we still lack a simple framework to explain what drives robustness across varying degrees of extremeness.
@@
To address this gap, we introduce the concept of super-resolution of extremal distributions, a novel approach that focuses on learning to increase the spatial resolution of the parameters governing precipitation extremes rather than reconstructing precipitation fields. This shift from traditional field-based super-resolution is motivated by tractability and theoretical consistency. Fine-resolution and coarse-resolution extremes do not occur simultaneously, but they typically fall within the same temporal block, making them only weakly paired. Such imperfect correspondence makes field-based super-resolution challenging, whereas distributional modeling naturally accommodates it by focusing on aggregated statistical properties rather than individual values. Moreover, super-resolution of distributions is generally not treated on a per-GCM basis. This contrasts with the full empirical downscaling pipeline required for local climate projections, where global and sometimes regional climate model outputs must be bias-corrected (francois2020multivariate; cannon2018multivariate; vrac2015multivariateBC) before being brought to the desired spatial scale. Beyond tractability, working with distributions offers several advantages. It aligns with practical needs for risk assessment at the local level: decision-makers often rely on quantiles or return levels derived from distributions, not raw precipitation fields. It may reduce the risk of overfitting to local patterns and enhances generalization to other regions and climates, as the model learns a small set of interpretable parameters rather than high-dimensional fields. Finally, this approach is computationally efficient, requiring fewer data and enabling interpretable models with lower training costs. friederichs2010statistical previously downscaled precipitation block maxima by targeting GEV parameters from large-scale circulation features. However, we focus here on cases where convection-permitting extremes (1–5 km) are at least partially resolved at coarser resolutions (12–50 km), consistent with the HighResMIP2 range targeted by next-generation of Earth system models (roberts2025highresmip2), allowing a super-resolution setup.
@@
Our generative super-resolution model combines regression-based modeling with extreme value theory (vrac2007stochastic). It learns the parameters of the Generalized Extreme-Value (GEV) distribution governing fine-resolution precipitation maxima from those governing their coarse-resolution counterparts. To ensure interpretability, we use Vector Generalized Additive Models (VGAM) and Vector Generalized Linear Models (VGLM), which allow us to specify a distribution family and examine how features influence predicted distributions through spline terms. This approach offers three key advantages: low training costs, reliable probability estimates, and transparent model behavior.
@@
To quantify our models’ robustness to climate change, we adopt a pseudo-reality framework over Switzerland, where complex topography—ranging from alpine areas above 4,000 m to low-lying regions below 200 m—drives strong spatial variability in precipitation. Observations show a positive trend in annual maximum daily precipitation at 91% of Swiss stations, with 10-minute summer precipitation intensities increasing at 5.7% per decade (bauer2024observedevolutionprecip; scherrer2016trends). Convection-permitting model ensembles reproduce these changes more faithfully than coarser models and project a 6-7% increase in heavy summer precipitation intensity per degree of warming, despite an overall summer mean decrease (estermann2025projections; ban2021first). As summer precipitation is dominated by convective events, it remains particularly challenging to downscale without underestimating variability (zubler2014localized; schmidli2007downscaling), though this can be partly mitigated by explicitly incorporating temperature dependence in precipitation scaling (moraga2024subdailyrainfall).
@@
Motivated by the challenge of super-resolving extreme summer precipitation, we base our pseudo-reality experiment on the present-future convection-permitting simulation pair of Hentgen2019data (Section 2). We then formalize the task of super-resolving distributions and define the “robustness gap” to quantitatively assess model generalization across climates (Section 3). Applied to the super-resolution of the GEV using VGAMs and VGLMs (Section 4), our diagnostic framework explains generalization errors across quantiles, interprets model behavior via splines, and identifies performance limits when the super-resolution factor becomes too large (Section 5). We conclude in Section 6. The Supplementary Material (SM) provides technical derivations that support Sections 4 and 5 of the manuscript.
@@ 2 Data
We super-resolve hourly precipitation model output over Switzerland for 11 European summer (JJA) seasons in both historical (1999–2009) and projected (2079–2089) climates. We focus on summer months to isolate convective precipitation; Swiss summers are characterized by weak synoptic forcing, frequent afternoon thunderstorms, and convective rainfall that are particularly challenging to downscale over the Alps. All data are derived from simulations using the regional weather and climate model COSMO (Consortium for Small-scale Modeling), run over a European domain of approximately $3000\times 3000$ km at 2.2 km grid spacing (Hentgen2019data). An overview of the simulations is given in Section 2.1; Leutwyler2017 documents prior validation of the control run but the associated model data and observations are not part of this study’s pipeline. Section 2.2 describes the construction of the coarse-resolution field from the fine-resolution data. The elevation statistics employed are detailed in Section 2.3, and the training, validation, and test set configurations are outlined in Section 2.4.
@@ 2.1 Simulations
We calibrate and test our method using two regional climate simulations at 2.2 km horizontal resolution from Hentgen2019data. Kilometer-scale climate modeling has become increasingly common in recent years (ban2021first; Stevens_2019), offering several advantages over coarser-resolution simulations. The finer representation of topography and land surface enables more realistic precipitation patterns in complex terrain such as the Alps. In addition, vertical air motion is explicitly resolved by the governing equations, bringing the model formulation closer to physical first principles. Compared to coarser resolutions that require convective parameterization, explicit convection improves the realism of the hydrological cycle, particularly for extreme precipitation and its associated mechanisms and feedbacks (Prein_2020; Schar_2020; Lenderink_2025).
To assess generalization, we use model output from a simulation based on the pseudo-global warming (PGW) method (Rasmussen_2011; Adachi_2020; Schar_1996). PGW aims to simulate a warmer climate by preserving the spatiotemporal structure of historical weather patterns. A regional simulation is first performed using reanalysis-based boundary conditions, referred to as the control simulation (CTRL). CTRL (1999–2008) was observationally validated in Leutwyler2017 against the ENSEMBLES E-OBS gridded European daily temperature and precipitation dataset (haylock2008EOBS), the EURO4M Alpine Precipitation Grid Dataset (APGD, isotta2014APGD), the Swiss radar-disaggregated hourly precipitation dataset RdisaggH (wuest2010RdisaggH), the German radar-disaggregated hourly precipitation dataset DisaggDE (paulat2008), solar surface radiation from the Baseline Surface Radiation Network (BSRN, hakuba2014), and lightning from the UK Met Office Arrival Time Difference network (ATDnet, anderson2014lightning).
A physically consistent climate change signal (“climate delta”) is then applied to the CTRL simulation’s boundaries, and the simulation is repeated. In Hentgen2019data, the deltas are 30-year climatological differences: 2070–2099 (RCP8.5) minus 1971–2000, taken from CMIP5 runs of the Max Planck Institute Earth System Model Low Resolution (MPI-ESM-LR) GCM (kroner2017separating). These deltas vary with latitude, longitude, elevation, and month. The PGW simulation hence retains the sequence of weather events from CTRL, adjusted by the imposed climate signal. As such, the PGW simulation primarily captures thermodynamic changes (e.g., increases in temperature and moisture) while preserving large-scale circulation patterns from the historical record (Hall_2024). This reduces confounding effects from circulation biases, providing a dataset of intermediate complexity with a clear RCP8.5 warming signal to stress-test the generalization of statistical downscaling methods.
@@ 2.2 Coarse-Graining
To ensure comparability across climates, we adopt an idealized super-resolution framework in which the coarse-resolution inputs are coarsened versions of the fine-resolution targets. Hourly precipitation data are aggregated to coarser grid spacings via mean pooling, i.e., using a square spatial filter of prescribed length on the native COSMO grid. To assess robustness across not only climates but also super-resolution factors, we apply pooling lengths of 13.2 km ($\times$6), 26.4 km ($\times$12), and 52.8 km ($\times$24). We then compute monthly maxima on the coarsened low-resolution data (features) and on the fine-resolution data (targets).
@@
As shown in Figure 1, decreasing spatial resolution leads to substantial information loss. Precipitation values become more uniform, with reduced means and variances of extremes consistent with the geostatistical concept of “change of support” (onibon2004; no-wook2023). The highest extremes are particularly affected, with distribution tails shortening as resolution decreases. This is due to the localized nature of extremes: coarse-resolution cells blend multiple events within a grid block, smoothing out their intensity and variability. These patterns remain consistent across all resolutions in both climates. Future climates exhibit broader, flatter distributions with wider supports, heavier tails and higher standard deviations, indicating increased dispersion.
@@ 2.4 Training, Validation, and Test Split
To prevent overfitting and to ensure objective model evaluation, we define spatially separated splits (valavi2019; brenning2012), shown in Figure 2. Rather than assigning entire contiguous regions to each split, we partition the domain into 10 spatial regions, which are then distributed across the training, validation, and test sets. This approach preserves spatial separation while maintaining representativeness of the overall data distribution in each set. The training set (white) includes 70% of the data and is used to optimize the model’s trainable parameters, providing enough information to learn the super-resolution mapping. The validation set (dark gray), comprising 17% of the data, is used for model and feature selection and helps prevent overfitting. The remaining 13% forms the test set (light gray), which evaluates the model’s ability to generalize to unseen regions. We refer to the generalization quantified by the test set as “spatial generalizability” to distinguish it from the models’ ability to generalize across climates.
@@ 3.1 Super-Resolution of Distributions
A concrete example for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of regular two-dimensional grids defined over the study domain $\mathcal{S}$, with mesh sizes $\delta_{1}>0$ and $\delta_{2}>0$.
$\mathcal{G}^{\delta_{1}}\overset{\mathrm{def}}{=}\left\{(\delta_{1}\cdot i_{1},\delta_{1}\cdot i_{2})\in\mathcal{S}:(i_{1},i_{2})\in\mathbb{N}^{2}\right\},$ (2)
$\mathcal{G}^{\delta_{2}}\overset{\mathrm{def}}{=}\left\{(\delta_{2}\cdot i_{1},\delta_{2}\cdot i_{2})\in\mathcal{S}:(i_{1},i_{2})\in\mathbb{N}^{2}\right\},$ (3)
where $\delta_{1}>\delta_{2}$, so that $\mathcal{G}^{\delta_{2}}$ has higher spatial resolution than $\mathcal{G}^{\delta_{1}}$.
@@
where $\bm{\tau}_{\bm{s}}$ represents a set of features specific to grid point $\bm{s}$ (e.g., latitude, longitude, topography, or land-use features), and $f$ is a function—possibly learned from data—that maps coarse-resolution distributional information and local covariates to fine-resolution distribution parameters.